# Moving Beyond Parental Control toward Community-based Approaches to Adolescent Online Safety

### v1

@@ 1. Introduction
This position paper examines empirical studies that explore this shift from parental control to collaborative parent-teen partnerships for establishing healthy online boundaries and jointly managing online safety, privacy, and security. Furthermore, we extend this collaborative model beyond the immediate family (Akter et al., 2024) to include broader community support networks (e.g., extended family, friends, neighbors, and colleagues) (Akter et al., 2023a, 2025, b). Finally, we pose open questions for the research community to consider during the workshop. Our position paper is highly relevant to the CHI 2025 Mobile Technology and Teens workshop as our approaches to adolescent online safety are one of the major themes of the workshop (i.e., engaging broader stakeholders and navigating the implications of emerging technologies).
@@ 3. From Parent-Teen Collaboration to Community-based Approach
To further explore the effectiveness of the collaborative approach to mobile online safety, security, and privacy, we recently conducted a field study examining how family members, including teens, collaborate with their broader trusted communities, not just extended relatives, such as friends, neighbors, and co-workers (Akter et al., 2025). From several prior studies, we learned that individuals with greater technological expertise often act as “caregivers,” providing support and guidance to those in the community who are less tech-savvy (”caregivees”) (Van Parys, 2019; Correa et al., 2015; Kiesler et al., 2000; Poole et al., 2009). We examined the distinct characteristics of these roles within communities and evaluated whether community-based approaches, such as Co-oPS, that treat individuals equally can address such caregiving disparities within communities. In our study, participants self-formed groups within their trusted networks and used Co-oPS for one month, engaging in collaborative digital safety and privacy management. Interestingly, we found that adolescent participants primarily assumed the role of caregivees, while adult participants acted as caregivers. This finding aligns with our other research (Akter et al., 2022c), which indicated that teens generally exhibit low interest in providing digital oversight for their parents. Furthermore, the study revealed that caregivees benefited significantly more from community-based oversight than caregivers. These results suggest that while teens may not take the initiative in promoting online safety, structured community-based interventions can effectively support their digital well-being. Overall, our findings underscore the potential for extending collaborative oversight beyond the family unit, fostering a broader, community-driven approach to adolescent digital safety and security.

### v3

@@ 1. Introduction
This position paper examines empirical studies that explore this shift from parental control to collaborative parent-teen partnerships for establishing healthy online boundaries and jointly managing online safety, privacy, and security. Furthermore, we extend this collaborative model beyond the immediate family (Akter et al., 2024a) to include broader community support networks (e.g., extended family, friends, neighbors, and colleagues) (Akter et al., 2023a, 2025, b, 2024b). Finally, we pose open questions for the research community to consider during the workshop. Our position paper is highly relevant to the CHI 2025 Mobile Technology and Teens workshop as our approaches to adolescent online safety are one of the major themes of the workshop (i.e., engaging broader stakeholders and navigating the implications of emerging technologies).
@@ 3. From Parent-Teen Collaboration to Community-based Approach
To further explore the effectiveness of the collaborative approach to mobile online safety, security, and privacy, we recently conducted a field study examining how family members, including teens, collaborate with their broader trusted communities, not just extended relatives, such as friends, neighbors, and co-workers (Akter et al., 2025). From several prior studies, we learned that individuals with greater technological expertise often act as “caregivers,” providing support and guidance to those in the community who are less tech-savvy (”caregivees”) (Van Parys, 2019; Correa et al., 2015; Kiesler et al., 2000; Poole et al., 2009; Kropczynski et al., 2021). We examined the distinct characteristics of these roles within communities and evaluated whether community-based approaches, such as Co-oPS, that treat individuals equally can address such caregiving disparities within communities. In our study, participants self-formed groups within their trusted networks and used Co-oPS for one month, engaging in collaborative digital safety and privacy management. Interestingly, we found that adolescent participants primarily assumed the role of caregivees, while adult participants acted as caregivers. This finding aligns with our other research (Akter et al., 2022c), which indicated that teens generally exhibit low interest in providing digital oversight for their parents. Furthermore, the study revealed that caregivees benefited significantly more from community-based oversight than caregivers. These results suggest that while teens may not take the initiative in promoting online safety, structured community-based interventions can effectively support their digital well-being. Overall, our findings underscore the potential for extending collaborative oversight beyond the family unit, fostering a broader, community-driven approach to adolescent digital safety and security.